OpenStudy (anonymous):

Set following equation up in partial fraction decomposition form:

OpenStudy (anonymous):

\[\frac { 5{ x }^{ 2 }-2x+3 }{ ({ x }^{ 2 }+1)(x-1) } \]

OpenStudy (anonymous):

\[5x^2-2x+3=\frac{ A }{ x^2+1 }+\frac{ B }{ x-1 }\]

OpenStudy (anonymous):

I need to correct something

OpenStudy (anonymous):

ya, just wanted to make sure this is the correct approach. thx.

OpenStudy (anonymous):

oh wait it should be Ax+B there instead of A i think

OpenStudy (anonymous):

The left should be the original equation, the right is correct. Did you understand that or would you like to see it?

OpenStudy (anonymous):

ya i understand that

OpenStudy (anonymous):

but i think you have to make the right \[\frac{Ax+B}{x^2+1}+\frac{B}{x-1}\]

OpenStudy (anonymous):

I think you are right except the next fraction should be a "C"

OpenStudy (anonymous):

In my notes I show that if you have (x^2 + c) you do what you did with Ax+B. (x+c) is just an A. But I show no repeated letters.

OpenStudy (anonymous):

yeah wow typo

OpenStudy (anonymous):

getting late

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

\[\frac{ 5x^2-2x+3 }{ (x^2+1)(x-1) }=\frac{ Ax+B }{ x^2+1 }+\frac{ C }{ x-1 }\]

OpenStudy (anonymous):

your welcome good luck